# MER 2025: When Affective Computing Meets Large Language Models

### v1

@@ Abstract.
MER2025 is the third year of our MER series of challenges, aiming to bring together researchers in the affective computing community to explore emerging trends and future directions in the field. Previously, MER2023111http://merchallenge.cn/mer2023 focused on multi-label learning, noise robustness, and semi-supervised learning, while MER2024222https://zeroqiaoba.github.io/MER2024-website introduced a new track dedicated to open-vocabulary emotion recognition. This year, MER2025 centers on the theme “When Affective Computing Meets Large Language Models (LLMs)”. We aim to shift the paradigm from traditional categorical frameworks reliant on predefined emotion taxonomies to LLM-driven generative methods, offering innovative solutions for more accurate and reliable emotion understanding. The challenge features four tracks: MER-SEMI focuses on fixed categorical emotion recognition enhanced by semi-supervised learning; MER-FG explores fine-grained emotions, expanding recognition from basic to nuanced emotional states; MER-DES incorporates multimodal cues (beyond emotion words) into predictions to enhance model interpretability; MER-PR investigates whether emotion prediction results can improve personality recognition performance. For the first three tracks, baseline code is available at MERTools [^3] https://github.com/zeroQiaoba/MERTools and datasets can be accessed via Hugging Face [^4] https://huggingface.co/datasets/MERChallenge/MER2025. For the last track, the dataset and baseline code are available on GitHub [^5] https://github.com/cai-cong/MER25_personality.
@@ 1. Introduction
Emotion plays a pivotal role in human-computer interactions (Picard, 2000; Zhou et al., 2018), attracting growing attention within the field of artificial intelligence. In the future, humans may interact more frequently with AI-powered systems than through direct human-to-human interactions. Emotion-aware interaction systems hold significant potential across various domains, including healthcare, banking, transportation, and education.
Human emotion expression is a complex process that typically involves multiple modalities, including facial expressions, vocal tones, body movements, gestures, and even physiological signals (Izard, 2009). This complexity has spurred the development of Multimodal Emotion Recognition (MER), a critical task that aims to integrate cross-modal cues to identify human emotions. Recently, MER research has evolved in two key directions: a) from coarse-grained (Lian et al., 2024a) to fine-grained emotion recognition (Lian et al., 2024b), and b) from categorical approaches (Lian et al., 2021) to descriptive emotion understanding (Lian et al., 2025), aiming to enhance both prediction accuracy and interpretability. In line with these advancements, MER2025@ACM Multimedia introduces four tracks aligned with current research priorities:
Track 1. MER-SEMI. Recent studies have shown that pre-training on large-scale unlabeled data, especially domain-matched data, can significantly enhance model performance (Lian et al., 2024a). This track provides a substantial collection of unlabeled samples from the same domain as the labeled data. Participants are encouraged to leverage semi-supervised learning techniques, such as masked auto-encoders (Tong et al., 2022) or contrastive learning (Radford et al., 2021), to achieve better results.
Track 2. MER-FG. Current frameworks primarily focus on basic emotions, often failing to capture the complexity and subtlety of human emotions. This track shifts the emphasis to fine-grained MER, enabling the prediction of a broader range of emotions. Following previous works (Lian et al., 2024b, 2025), participants are encouraged to leverage large language models (LLMs) for this purpose. Given that LLMs possess extensive vocabularies, they hold the potential to generate more diverse emotion categories beyond basic labels.
@@
Figure 1 shows the differences between MER-SEMI, MER-FG, and MER-DES. MER2025 builds upon the success of MER 2023@ACM Multimedia (Lian et al., 2023a) and MER 2024@IJCAI (Lian et al., 2024c). Over the course of these challenges, participation has steadily grown, increasing from 76 teams in MER2023 to 94 teams in MER2024. This year, we aim to attract more teams to join our challenge. Through the MER series of challenges, we strive to establish a unified platform for comparing different systems and further advancing the development of MER.
Figure 1. Demo for MER-SEMI, MER-FG, and MER-DES.
@@ 3.2. Evaluation Metrics
This track does not restrict the number or category of labels during prediction. Therefore, traditional metrics like WAF are not suitable for MER-FG. For evaluation purposes, we follow the previous work (Lian et al., 2024b) and compute results in two stages.
@@ 4.2. Evaluation Metrics
MER-DES uses multimodal cues to enhance the interpretability of each emotion label. Since interpretability is a subjective metric, we plan to recruit annotators to manually score the emotion descriptions. The scoring will be conducted along two dimensions:
@@
For each team, we will request that two members assist with the scoring process. The manual annotations will then be used to determine the final rankings.
@@ 4.3.1. Zero-shot Baselines
We can also use MLLMs as baselines. Specifically, we can prompt the MLLMs with the following instructions: As an expert in the field of emotions, please focus on the facial expressions, body movements, environment, acoustic information, subtitle content, etc., in the video to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the video.
@@ 4.3.2. AffectGPT
We can also fine-tune AffectGPT using the emotion descriptions from the OV-MERD and MER-Caption+ datasets. The official GitHub repository already includes baseline code and pretrained weights for this track.
5. MER-PR
5.1. Dataset
This track aims to explore whether emotion prediction results can enhance the performance of personality recognition. For this purpose, we utilize the MDPE dataset (Cai et al., 2024), which provides annotations for both emotion and personality traits. The dataset and baseline code are provided by Xuefei Liu and are available on GitHub [^6] https://github.com/cai-cong/MER25_personality. Table 5 provides the dataset statistics.
For personality traits, we used a Big Five personality questionnaire consisting of 60 items. Responses were used to derive scores for five personality dimensions (extraversion, agreeableness, conscientiousness, neuroticism, and openness), with each dimension represented as a floating-point value ranging from 0 to 1. The dataset includes 233 participants, all native Chinese speakers from diverse backgrounds. Participants ranged in age from 18 to 69 years and held various professions.
@@ 5.2. Evaluation Metric
Table 6. Unimodal results for MER-PR. In this table, the abbreviations “E”, “A”, “C”, “N”, and “O” represent the extraversion, agreeableness, conscientiousness, neuroticism, and openness, respectively.
Feature Val Test
E()($\downarrow$)( ↓ ) A ()($\downarrow$)( ↓ ) C ()($\downarrow$)( ↓ ) N ()($\downarrow$)( ↓ ) O ()($\downarrow$)( ↓ ) Avg ()($\downarrow$)( ↓ ) E ()($\downarrow$)( ↓ ) A ()($\downarrow$)( ↓ ) C ()($\downarrow$)( ↓ ) N ()($\downarrow$)( ↓ ) O()($\downarrow$)( ↓ ) Avg ()($\downarrow$)( ↓ )
VIT 0.174 0.231 0.159 0.183 0.179 0.185 0.178 0.217 0.106 0.156 0.145 0.160
ClipVIT-B16 0.177 0.225 0.169 0.187 0.171 0.185 0.180 0.220 0.108 0.165 0.158 0.166
ClipVIT-L14 0.180 0.230 0.168 0.180 0.172 0.186 0.176 0.229 0.114 0.166 0.173 0.172
HUBERT-base 0.176 0.228 0.162 0.184 0.173 0.185 0.183 0.218 0.098 0.167 0.160 0.165
HUBERT-large 0.194 0.240 0.171 0.200 0.200 0.201 0.210 0.228 0.150 0.174 0.189 0.190
Wav2vec2-base 0.172 0.221 0.160 0.182 0.169 0.180 0.177 0.215 0.092 0.158 0.159 0.160
Wav2vec2-large 0.190 0.237 0.168 0.200 0.198 0.198 0.205 0.223 0.146 0.173 0.190 0.187
WavLM-base 0.179 0.227 0.164 0.184 0.178 0.186 0.180 0.219 0.101 0.162 0.156 0.164
WavLM-large 0.191 0.246 0.174 0.206 0.191 0.201 0.209 0.251 0.147 0.205 0.200 0.202
Sentence-BERT 0.176 0.234 0.181 0.181 0.181 0.190 0.181 0.219 0.104 0.163 0.159 0.165
ChatGLM2-6B 0.195 0.248 0.180 0.202 0.188 0.203 0.200 0.240 0.133 0.180 0.182 0.187
Baichuan-13B 0.172 0.229 0.160 0.180 0.167 0.182 0.157 0.214 0.098 0.157 0.156 0.156
Table 7. Multimodal results for MER-PR.
Features Val Test
V A T E()($\downarrow$)( ↓ ) A ()($\downarrow$)( ↓ ) C ()($\downarrow$)( ↓ ) N ()($\downarrow$)( ↓ ) O ()($\downarrow$)( ↓ ) Avg ()($\downarrow$)( ↓ ) E ()($\downarrow$)( ↓ ) A ()($\downarrow$)( ↓ ) C ()($\downarrow$)( ↓ ) N ()($\downarrow$)( ↓ ) O()($\downarrow$)( ↓ ) Avg ()($\downarrow$)( ↓ )
VIT W2V 0.175 0.223 0.160 0.185 0.171 0.183 0.177 0.213 0.095 0.158 0.159 0.160
W2V BAI 0.171 0.223 0.160 0.184 0.170 0.182 0.175 0.215 0.096 0.160 0.159 0.161
VIT BAI 0.172 0.228 0.161 0.182 0.167 0.182 0.175 0.216 0.097 0.158 0.154 0.160
VIT W2V BAI 0.175 0.224 0.160 0.187 0.172 0.184 0.171 0.212 0.093 0.160 0.156 0.158
@@ 6. Challenge Protocol
For MER-SEMI, MER-FG, and MER-DES, to download the dataset, participants must complete an End User License Agreement (EULA) available on Hugging Face [^7] https://huggingface.co/datasets/MERChallenge/MER2025. The EULA clearly states that the dataset is for academic research purposes only and prohibits any modifications or uploads to the Internet. We will review the submitted EULA promptly; once approved, participants will gain access to the dataset. We emphasize that manual annotation of MER2025 samples is strictly prohibited. For additional requirements, please visit our official website [^8] https://zeroqiaoba.github.io/MER2025-website. For these tracks, the test samples are selected from the 124k unlabeled samples (see Table 1). To reduce the task difficulty, we reduce the evaluation scope from 124k to 20k samples. Participants must submit predictions for these 20k samples, which cover all the test samples of these tracks. As shown in Figure 1, each track has distinct objectives. For MER-SEMI, participants must predict the most likely label from six predefined categories: worried, happy, neutral, angry, surprised, and sad; For MER-FG, participants can freely predict any emotion labels without restrictions on category or quantity; For MER-DES, participants are required to submit both multimodal evidence and corresponding emotion labels to improve model interpretability.
@@ 7. Conclusions
This year’s MER2025 centers on the theme “when affective computing meets large language models” and features four distinct tracks. MER-SEMI focuses on categorical emotion recognition using a fixed taxonomy, leveraging semi-supervised learning and unlabeled samples to enhance model performance. MER-FG specializes in fine-grained emotion recognition, allowing flexible label numbers and emotion categories during prediction without constraints. MER-DES extends beyond emotion label prediction by requiring participants to provide supporting multimodal clues for their predictions. MER-PR explores the interplay between emotion and personality, investigating whether emotion labels can improve model performance in personality prediction. This paper outlines the datasets, baselines, evaluation metrics, and experimental results for all four tracks. To facilitate reproducibility, we have open-sourced the code and pretrained weights on our official GitHub repository. Through the MER series challenges, we aim to unite researchers in this field and collectively advance the development of affective computing.

### v2

@@ Abstract.
MER2025 is the third year of our MER series of challenges, aiming to bring together researchers in the affective computing community to explore emerging trends and future directions in the field. Previously, MER2023111http://merchallenge.cn/mer2023 focused on multi-label learning, noise robustness, and semi-supervised learning, while MER2024222https://zeroqiaoba.github.io/MER2024-website introduced a new track dedicated to open-vocabulary emotion recognition. This year, MER2025 centers on the theme “When Affective Computing Meets Large Language Models (LLMs)”. We aim to shift the paradigm from traditional categorical frameworks reliant on predefined emotion taxonomies to LLM-driven generative methods, offering innovative solutions for more accurate and reliable emotion understanding. The challenge contains four tracks: MER-SEMI focuses on fixed categorical emotion recognition enhanced by semi-supervised learning; MER-FG explores fine-grained emotions, expanding recognition from basic to nuanced emotional states; MER-DES incorporates multimodal cues (beyond emotion words) into predictions to enhance model interpretability; MER-PR reveals whether emotion prediction results can improve personality recognition performance. For the first three tracks, the baseline code is available at MERTools [^3] https://github.com/zeroQiaoba/MERTools and datasets can be accessed via Hugging Face [^4] https://huggingface.co/datasets/MERChallenge/MER2025. For the last track, the dataset and baseline code are available on GitHub [^5] https://github.com/cai-cong/MER25_personality.
@@ 1. Introduction
Emotion plays a pivotal role in human-computer interactions (Picard, 2000; Zhou et al., 2018), drawing increasing attention within the field of artificial intelligence. Already today, humans interact more frequently with AI systems than through direct human-to-human interaction, and this trend is expected to grow even more pronounced in the coming years. Emotion-aware interaction systems hold significant potential across various domains, including healthcare, banking, transportation, and education (Feidakis, 2016; Haddad et al., 2024).
Human emotion expression is a complex process that typically involves multiple modalities, including facial expressions, vocal tones, body movements, gestures, and even physiological signals (Izard, 2009). This complexity has spurred the development of Multimodal Emotion Recognition (MER), a critical task that aims to integrate cross-modal cues to identify human emotions. Recently, MER research has evolved in two key directions: a) from coarse-grained (Lian et al., 2024a) to fine-grained emotion recognition (Lian et al., 2024b), and b) from categorical approaches (Lian et al., 2021) to descriptive emotion understanding (Lian et al., 2025; Cheng et al., 2024), aiming to enhance both prediction accuracy and interpretability. In line with these advancements, MER2025@ACM Multimedia introduces four tracks aligned with current research priorities:
Track 1. MER-SEMI. Recent studies have demonstrated that pre-training on large-scale unlabeled data (Ma et al., 2024), especially domain-matched data, can significantly enhance model performance (Lian et al., 2024a; Sun et al., 2023; Cheng et al., 2023). This track provides a substantial collection of unlabeled samples from the same domain as the labeled data. Participants are encouraged to leverage semi-supervised learning techniques, such as masked auto-encoders (Tong et al., 2022) or contrastive learning (Radford et al., 2021), to achieve better results.
Track 2. MER-FG. Current frameworks primarily focus on basic emotions, often failing to capture the complexity and subtlety of human emotions. This track shifts the emphasis to fine-grained MER, enabling the prediction of a broader range of emotions. Following previous works (Lian et al., 2024b; Lian et al., 2025), participants are encouraged to leverage large language models (LLMs) for this purpose. Given that LLMs possess extensive vocabularies, they hold the potential to generate more diverse emotion categories beyond basic labels.
@@
Figure 1 shows the differences between these tracks. MER2025 builds upon the success of MER 2023@ACM Multimedia (Lian et al., 2023a) and MER 2024@IJCAI (Lian et al., 2024c). Over the course of these challenges, participation has steadily grown, increasing from 76 teams in MER2023 to 94 teams in MER2024. This year, we aim to attract more teams to join our challenge. Through the MER series of challenges, we strive to establish a unified platform for comparing different systems and further advancing the development of MER.
Figure 1. Demo for four tracks.
@@ 3.2. Evaluation Metrics
This track does not restrict the number or category of labels during prediction. Therefore, traditional metrics like accuracy are not suitable for MER-FG. For evaluation purposes, we follow the previous work (Lian et al., 2024b) and compute results in two stages.
@@ 4.2. Evaluation Metrics
MER-DES leverages multimodal cues to enhance the interpretability of each emotion label. Given that interpretability is inherently a relatively subjective metric, we plan to recruit annotators to manually evaluate the quality of the emotion descriptions. Simultaneously, we will invite two members from each team to assist with the scoring process. The evaluation will be conducted along two dimensions:
@@
Initially, we plan to provide real descriptions and calculate similarity scores between real and predicted descriptions, as done in EMER (Lian et al., 2023b). However, human emotional expression is complex and difficult to fully capture all emotion-related visual and acoustic cues. Therefore, we shift our focus to manual evaluation. During the ranking process, each team is required to participate in the labeling process. We will exclude annotators with low inter-annotator agreement compared to others to ensure annotation quality. Additionally, we will mask the sample names and require participants to submit results in English while using their translated Chinese versions for ranking. This approach prevents participants from identifying their own submissions and deliberately assigning inflated scores. The final rankings will be determined based on these manual annotations.
@@ 4.3.1. Zero-shot Baselines
You can also use MLLMs as baselines. Specifically, you can prompt the MLLMs with the following instructions: As an expert in the field of emotions, please focus on the facial expressions, body movements, environment, acoustic information, subtitle content, etc., in the video to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the video.
@@ 4.3.2. AffectGPT
You can also fine-tune AffectGPT using the emotion descriptions from the OV-MERD and MER-Caption+ datasets. The official GitHub repository already includes baseline code and pretrained weights for this track.
@@
Table 6. Unimodal results for MER-PR. In this table, the abbreviations “O”, “C”, “E”, “A”, and “N” represent the openness, conscientiousness, extraversion, agreeableness, and neuroticism, respectively.
Feature Val Test
O()($\downarrow$)( ↓ ) C ()($\downarrow$)( ↓ ) E ()($\downarrow$)( ↓ ) A ()($\downarrow$)( ↓ ) N ()($\downarrow$)( ↓ ) Avg ()($\downarrow$)( ↓ ) O ()($\downarrow$)( ↓ ) C ()($\downarrow$)( ↓ ) E ()($\downarrow$)( ↓ ) A ()($\downarrow$)( ↓ ) N()($\downarrow$)( ↓ ) Avg ()($\downarrow$)( ↓ )
VIT 0.179 0.159 0.174 0.231 0.183 0.185 0.145 0.106 0.178 0.217 0.156 0.160
ClipVIT-B16 0.171 0.169 0.177 0.225 0.187 0.185 0.158 0.108 0.180 0.220 0.165 0.166
ClipVIT-L14 0.172 0.168 0.180 0.230 0.180 0.186 0.173 0.114 0.176 0.229 0.166 0.172
HUBERT-base 0.173 0.162 0.176 0.228 0.184 0.185 0.160 0.098 0.183 0.218 0.167 0.165
HUBERT-large 0.200 0.171 0.194 0.240 0.200 0.201 0.189 0.150 0.210 0.228 0.174 0.190
Wav2vec2-base 0.169 0.160 0.172 0.221 0.182 0.180 0.159 0.092 0.177 0.215 0.158 0.160
Wav2vec2-large 0.198 0.168 0.190 0.237 0.200 0.198 0.190 0.146 0.205 0.223 0.173 0.187
WavLM-base 0.178 0.164 0.179 0.227 0.184 0.186 0.156 0.101 0.180 0.219 0.162 0.164
WavLM-large 0.191 0.174 0.191 0.246 0.206 0.201 0.200 0.147 0.209 0.251 0.205 0.202
Sentence-BERT 0.181 0.181 0.176 0.234 0.181 0.190 0.159 0.104 0.181 0.219 0.163 0.165
ChatGLM2-6B 0.188 0.180 0.195 0.248 0.202 0.203 0.182 0.133 0.200 0.240 0.180 0.187
Baichuan-13B 0.167 0.160 0.172 0.229 0.180 0.182 0.156 0.098 0.157 0.214 0.157 0.156
Table 7. Multimodal results for MER-PR.
Features Val Test
V A T O()($\downarrow$)( ↓ ) C ()($\downarrow$)( ↓ ) E ()($\downarrow$)( ↓ ) A ()($\downarrow$)( ↓ ) N ()($\downarrow$)( ↓ ) Avg ()($\downarrow$)( ↓ ) O ()($\downarrow$)( ↓ ) C ()($\downarrow$)( ↓ ) E ()($\downarrow$)( ↓ ) A ()($\downarrow$)( ↓ ) N()($\downarrow$)( ↓ ) Avg ()($\downarrow$)( ↓ )
VIT W2V 0.171 0.160 0.175 0.223 0.185 0.183 0.159 0.095 0.177 0.213 0.158 0.160
W2V BAI 0.170 0.160 0.171 0.223 0.184 0.182 0.159 0.096 0.175 0.215 0.160 0.161
VIT BAI 0.167 0.161 0.172 0.228 0.182 0.182 0.154 0.097 0.175 0.216 0.158 0.160
VIT W2V BAI 0.172 0.160 0.175 0.224 0.187 0.184 0.156 0.093 0.171 0.212 0.160 0.158
5. MER-PR
5.1. Dataset
This track aims to explore whether emotion prediction results can enhance the performance of personality recognition. For this purpose, we utilize the MDPE dataset (Cai et al., 2024), which provides annotations for both emotion and personality traits. The dataset and baseline code are available on GitHub [^6] https://github.com/cai-cong/MER25_personality. Table 5 provides dataset statistics.
For personality traits, we used a Big Five personality questionnaire consisting of 60 items. Responses were used to derive scores for five personality dimensions (openness, conscientiousness, extraversion, agreeableness, and neuroticism), with each dimension represented as a floating-point value ranging from 0 to 1. The dataset includes 233 participants, all native Chinese speakers from diverse backgrounds.
@@ 6. Challenge Protocol
For MER-SEMI, MER-FG, and MER-DES, to download the dataset, participants must complete an End User License Agreement (EULA) available on Hugging Face [^7] https://huggingface.co/datasets/MERChallenge/MER2025. The EULA clearly states that the dataset is for academic research purposes only and prohibits any modifications or uploads to the Internet. We emphasize that manual annotation of MER2025 samples is strictly prohibited. For these tracks, the test samples are selected from the 124k unlabeled samples (see Table 1). To reduce the task difficulty, we reduce the evaluation scope from 124k to 20k samples. Participants must submit predictions for these 20k samples, which cover all the test samples of these tracks. As shown in Figure 1, each track has distinct objectives. For MER-SEMI, participants must predict the most likely label from six predefined categories: worried, happy, neutral, angry, surprised, and sad; For MER-FG, participants can freely predict any emotion labels without restrictions on category or quantity; For MER-DES, participants are required to submit both multimodal evidence and corresponding emotion labels to improve model interpretability.
@@ 7. Conclusions
This year’s MER2025 focuses on the theme “When Affective Computing Meets Large Language Models” and contains four distinct tracks. This paper presents the datasets, baselines, evaluation metrics, and experimental results for all tracks. For MER-SEMI, we evaluate various unimodal features and multimodal fusion approaches, providing a strong baseline for categorical emotion recognition under a fixed taxonomy. For MER-FG and MER-DES, we employ LLM-driven baselines to generate fine-grained emotions with corresponding multimodal clues, ensuring accurate and interpretable emotion understanding. For MER-PR, we test diverse unimodal features for personality detection. To promote reproducibility, we have open-sourced the code and pretrained models on our official GitHub repository. We hope all participants enjoy this year’s challenge! Your continued support and engagement make this challenge truly meaningful.